40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 12 00:14:00 2013, yuta, Update, Locking, PRM coil balancing 
    Reply  Sat Feb 16 17:27:26 2013, yuta, Update, SUS, PRM coil balanced PRM_f2a.png
       Reply  Sun Feb 17 19:27:19 2013, rana, Update, SUS, PRM coil balanced 
       Reply  Wed Feb 27 11:59:43 2013, yuta, Update, SUS, PRM/BS coil re-balanced PRM_f2a20130227.pngBS_f2a20130227.png
Message ID: 8065     Entry time: Tue Feb 12 00:14:00 2013     Reply to this: 8093
Author: yuta 
Type: Update 
Category: Locking 
Subject: PRM coil balancing 

We tried to lock half-PRC tonight, but we couldn't. Why?? I could lock yesterday.
It locks for ~ 1 sec, but it beam spot motion freaks out mainly in yaw.
I tried to balance PRM coils, but oplev beam was clipped by MMT1......

What I did:
  1. Found elog #5392 and found F2P_LOCKIN.py

  2. Modified F2P_LOCKIN.py because LOCKIN channel names are some how changed like this;

LOCKIN1_I -> LOCKIN1_DEMOD_I
LOCKIN1_Q -> LOCKIN1_DEMOD_Q
LOCKIN1_SIG -> LOCKIN1_DEMOD_SIG

  3. Running

/opt/rtcds/caltech/c1/scripts/SUS/F2P_LOCKIN.py -o PRM

  should adjust (UL|UR|LR|LL)COIL_GAINs by putting some gain imbalance and shaking the mirror in different frequencies. It uses LOCKIN to OL(PIT|YAW).

  4. Since there was no PRM oplev beam coming out from the vacuum, I quickly looked into BS-PRM chamber. Oplev beam was clipped by MMT1. If I adjust PRM slider values to avoid clipping, the beam will be clipped by mirrors on oplev table. What happened to the PRM oplev?

  5. I also made bunch of /opt/rtcds/userapps/trunk/sus/c1/medm/templates/SUS_SINGLE_LOCKIN(1|2)_DEMOD_(I|Q|SIG).adl because there were missing screens.

Next:
 We need to restore the PRM oplev and balance the coils. See, also, elog #7679

ELOG V3.1.3-